# Uncertainty, Single Crossing Property, and Stochastic Choice Data

### v1

@@
Uncertainty, Single Crossing Property, and Stochastic Choice Data
Tanay Raj Bhatt111trbhatt@utexas.edu, Department of Economics, University of Texas at Austin. I thank Abhishek Arora, V Bhaskar, Svetlana Boyarchenko, Francesco Conti, Faheem Gilani, Maxwell Stinchcombe, Vasiliki Skreta, Caroline Thomas, Tom Wiseman, and Leeat Yariv for helpful comments. Laurens Cherchye, Bram De Rock, and Thomas Demuynck provided crucial help and guidance during my research visit to Universite Libré de Bruxelles, where part of this work was undertaken. Finally, I thank seminar and conference audience at UT Austin, ULB, and the Delhi School of Economics for helpful discussion and comments. I also acknowledge the financial support received from Fonds de la Recherche Scientifique (FNRS), Belgium, which partly supported my research visit to ULB. Any remaining errors are my own.
(March 2025)
In a typical model of private information and choice under uncertainty, a decision maker observes a signal, updates her prior beliefs using Bayes rule, and maximizes her expected utility. If the decision maker’s utility function satisfies the single crossing property, and the information structure is ordered according to the monotone likelihood ratio, then the comparative statics exhibit monotonicity with respect to signals. We consider the restrictions placed by this model of signal processing on state conditional stochastic choice data. In particular, we show that this model rationalizes a state conditional stochastic choice dataset if and only if the dataset itself is ordered according to the monotone likelihood ratio.
@@
The single crossing property is an ordinal generalization of the increasing differences property [^3] A function $u$ exhibits increasing differences in $(x,\theta)$ if, $x^{\prime}>x$ and $\theta^{\prime}>\theta$ imply $u(x^{\prime},\theta^{\prime})-u(x,\theta^{\prime})>u(x^{\prime},\theta)-u(x,\theta)$. Clearly, this is a cardinal property. See Topkis, 1998.. Since it implies monotone comparative statics, and due to its straightforward economic interpretation resting on strategic complementarity, the SCP has turned out to be an extremely important tool for applications - both in applied theory as well as in the empirical literature. For instance, in mechanism design, single crossing property facilitates closed form solution for optimal transfers. In applied work, the latter has been used to solve certain optimal taxation problems (see Mirrlees, 1971). In IO models of entry, single crossing property is often invoked and facilitates equilibrium analysis. In labor economics, the single crossing property is closely related to the so called Le Chatelier principle which helps characterize comparative static results (see Dekel, Quah and Sinander, 2022). The single crossing property also has some nice implications for preference aggregation, which proves to be quite useful in applied work (see, for instance Gans and Smart, 1996).
@@
In this paper, we ask what (if any) restrictions does the single crossing property impose on certain kinds of observable data. Often, available data may take a stochastic form. The analyst may observe only some distribution(s) over available actions and may not possess any knowledge of either the DM’s utility function or her private information. It is therefore important to understand whether the behavior of a Bayes rational agent imposes any restrictions on the stochastic data observed by the analyst. Caplin and Martin, 2015 show that rationalizability by a Bayes rational agent with access to some private information is equivalent to a single condition. Here, we ask what if the single crossing and MLR-ordered information structure impose any additional restrictions on top of those identified by Caplin and Martin, 2015. The importance of this question is illustrated by the aforementioned examples. Since the SCP has proven to be such an important property, it seems worthwhile to see if it imposes meaningful testable restrictions on observable data.
@@
Athey’s results suggest that the stochastic data should exhibit certain kind of monotonicity. If $(i)$ the DM’s utility function is single crossing; and $(ii)$ private information structure is MLR-ordered, then we know from Athey, 2002 that $(iii)$ the comparative statics are monotone. How does $(i)-(iii)$ affect the data? In our setup, neither the utility function nor the private information structure is known, but only a set of conditional distributions over actions are observed. If $(ii)$ holds, then higher states lead to higher signal realizations more often - or with higher likelihood. If $(i)$ holds as well, then $(iii)$ is true, and so higher signals would lead directly to higher actions. One can expect this to be exhibited in data: since higher signals are more likely, and they lead to higher actions, we would expect higher actions to be more likely at higher states. This monotonicity translates into the data itself being MLR-ordered - that is, higher actions have higher relative probabilities at higher states. Proposition 12 shows that this intuition holds for the binary case. In fact, we establish something much stronger for the binary case. We show that this condition is not only both necessary and sufficient for rationalization by an SCP utility function, but such data can only be rationalized by a utility function satisfying the single crossing property. Furthermore, the main result of the paper shows that for any finite number of states, the MLR condition is both necessary and sufficient for rationalizability by a single crossing utility and an MLR ordered information structure.
@@
Related literature. Our paper tries to answer a question similar to ones asked within revealed preference theory, although the dataset we consider is quite different from finite data typically addressed in classical revealed preference theory. Green and Osband, 1991, Green and Srivastava, 1986, Echenique and Saito, 2015, Bayer et al., 2013 are prominent examples of classical revealed preference theory for expected utility. However, this set of papers considers a much richer setting. They focus primarily on consumption or asset demand, while none explicitly feature information processing or consider the single crossing property. Caplin and Martin, 2015, Doval and Eilat, 2023, Rehbeck, 2023 consider stochastic datasets and consider rationalizability by a Bayesian DM. The latter two consider rationalizability when the analyst only observes marginal distribution(s) over actions, but knows the DM’s utility function. Our paper can thus be thought of as imposing particular shape restrictions in the setting of Caplin and Martin, 2015. Lazzati, Quah and Shirai, 2018 consider similar shape restrictions, but they prove a rationalizability theorem with finite data, where the variation in data comes from “budgets” of actions. With finite data, they show that single crossing is indistinguishable from a weaker property developed by Quah and Strulovici, 2009, called interval dominance property. They also derive similar results for cross-sectional datasets and Bayesian games.
@@
Apesteguia, Ballester and Lu, 2017 posits a single crossing random utility model, dubbed SCRUM. The primary difference between that paper and ours is contextual: they deal with stochastic choice - there is a set of preferences with a probability measure defined on them that needs to be recovered. Here, the preferences are not stochastic - only the data is, and one is interested in the existence of some preference and information structure consistent with the given data. Put another way, while the structure of the dataset available to the analyst is somewhat similar across these two groups of papers, the postulated data generating process is different.
@@
The next section presents the model, characterizes the dataset available to the analyst, and defines the corresponding notion of rationalizability. Section 2 presents the main result for the binary case. The following two sections establish necessity and sufficiency, respectively, for the general case. Section 5 comprises of concluding discussion and outlines planned future work.
@@ 1.1 Preliminaries
Let $\Theta$ be a finite set of states and let $A$ be a finite set of actions, with $|A|\geq|\Theta|$. Typical elements of these sets are denoted by $\theta$ and $a$, respectively. Suppose that both $\Theta$ and $A$ are (strictly) ordered according to some exogenously given binary relations $>_{\Theta},>_{A}$. [^4] In particular, $>_{\Theta},>_{A}$ are assumed to be strict linear orders. That is, they are complete, transitive, and contain no “equalities”. While the results on which this paper builds are given for partial orders, we restrict attention to linear orders for brevity. Presence of unordered alternatives/states/signals does not affect the results reported here. To avoid notational clutter, in what follows, the subscripts are suppressed and the symbol $>$ is used to denote both the orderings. A decision maker (DM) wants to choose an action $a\in A$ without observing the state so as to maximize her the expectation of her utility function $u:A\times\Theta\rightarrow\mathbb{R}$ over states.
Let $S$ be some set of signal realizations, with $|S|\geq|\Theta|$. Let $S$ be ordered by some fixed binary relation $>_{S}$ as well. An information structure is a collection $\mathcal{I}\equiv\{\mu(\cdot|\theta)\}_{\theta\in\Theta}\subset\Delta(S)$. That is, each state induces a distribution over signal realizations. Let $\mu_{0}$ be the prior distribution over $\Theta$ [^5] To avoid unnecessary complications in the exposition, we assume that $\mu_{0}(\theta)>0$ for each $\theta\in\Theta$. This is without loss of generality since, if any state $\theta$ has a 0 probability, we can just drop that state and restrict attention to the set of states $\Theta/\{\theta\}$ instead. For revealed preference exercises, any such state is irrelevant.. The decision maker is said to be Bayesian if she obtains a signal realization $s\in S$, forms a belief $\mu(\cdot|s)\in\Delta(\Theta)$ using Bayes rule, and chooses an action $a^{*}(s)\in\arg\max_{A}\mathbb{E}_{\mu(\cdot|s)}[u(a,\theta)]$.
@@ 1.2 Data and rationalizability
Note, however, that it is possible that the DM optimally mixes between different actions. To that end, rationalizing the observed data also requires a choice rule $C:\Delta(\Theta)\rightarrow\Delta(A)$, mapping posterior beliefs into distributions over actions. Let $C(a|\gamma)$ denote the probability on action $a$ placed by the choice rule when the induced posterior is $\gamma$. Thus, an MBEU-DM is characterized by a tuple $\langle u,\mathcal{I},C\rangle$, where $u$ is a strict single crossing utility function, $\mathcal{I}$ is an MLR-ordered information structure, and $C$ is a choice rule. We say that the data $\mathcal{D}$ is rationalized by an MBEU-DM - represented by $\langle u,\mathcal{I},C\rangle$ if the distribution over actions induced by the optimal behavior of the DM characterized by the tuple is exactly $\bm{q}$. The following definition makes this notion of rationalizability precise,
@@
    Optimality: For all $\gamma\in$ supp$(\pi)$, and $a\in A$ such that $C(a|\gamma)>0$,
@@
The third condition demands that the observed data must in fact be induced by the choice behavior of the DM, while the fourth condition demands that the choice behavior be optimal.
2 Binary-state case
In the two-states case, the characterization turns out to be quite straightforward. Let there be two states and two actions so that $\Theta=\{\theta_{1},\theta_{2}\}$, $A=\{a_{1},a_{2}\}$, $a_{1}<a_{2}$ and $\theta_{1}<\theta_{2}$, and let the prior be $\mu_{0}=Pr(\theta=\theta_{1})$. The utility function is characterized by four numbers $u(a_{i},\theta_{j}),i,j=1,2$. The SCP is violated in this case if (and only if) $u(a_{1},\theta_{1})\leq u(a_{2},\theta_{1})$ and $u(a_{1},\theta_{2})>u(a_{2},\theta_{2})$. Denote the points in the dataset $q(a_{i}|\theta_{j})$ as $q^{i}_{j}$ so that subscripts correspond to states and the superscripts correspond to actions. Similarly, denote the utility $u(a_{i},\theta_{j})$ by $u^{i}_{j}$. Note that any dataset in this case is characterized by two numbers: $q(a_{1}|\theta_{1})\equiv q^{1}_{1},q(a_{1}|\theta_{2})\equiv q^{1}_{2}$. From these, we can compute $q^{2}_{1}=1-q^{1}_{1}$ and $q^{2}_{2}=1-q_{2}^{1}$. The MBEU hypothesis then has a straightforward test as demonstrated in the following lemma.
@@ 2 Binary-state case
Thus, in the binary case, the MLR property alone is both necessary and sufficient for there to exist an MBEU-rationalization. The above result suggests a straightforward test of the SCP. Indeed, if the data is not MLR-ordered, then no BEU-rationalization of the data can satisfy single crossing property. On the contrary, if the data is MLR-ordered, then there exists a BEU-rationalization that satisfies the SCP. The proof also exhibits something much stronger. It turns out that, in the binary case, an MLR-ordered dataset can be rationalized only by an MBEU-DM. That is, if there is an information structure and a utility function that rationalizes the data, then the Information structure must be MLR-ordered and utility function must be single crossing. This result is stated below as a corollary.
@@
From an empirical perspective, this result is crucial, since many econometric application of single crossing property involve binary states (for instance, entry games). The next sections establish that Theorem 1 holds in the general case as well.

### v3

@@
Testing Single Crossing Property with Stochastic Choice Data [^1] I thank Abhishek Arora, V Bhaskar, Svetlana Boyarchenko, Francesco Conti, Faheem Gilani, Raghav Malhotra, Maxwell Stinchcombe, Vasiliki Skreta, Caroline Thomas, Thomas Wiseman, and Leeat Yariv for helpful comments. Laurens Cherchye, Bram De Rock, and Thomas Demuynck provided crucial help and guidance during my research visit to Universite Libré de Bruxelles, where part of this work was undertaken. Sydney Neal provided excellent research assistance. Finally, I thank seminar and conference audiences at UT Austin, ULB, Delhi School of Economics, and Stony Brook University for helpful discussion and comments. I also acknowledge the financial support received from Fonds de la Recherche Scientifique (FNRS), Belgium, which partly supported my research visit to ULB. Any remaining errors are my own.
Tanay Raj Bhatt222trbhatt@utexas.edu, Department of Economics, University of Texas at Austin.
(September 2025)
In a typical model of private information and choice under uncertainty, a decision maker observes a signal, updates her prior beliefs using Bayes rule, and maximizes her expected utility. If the decision maker’s utility function satisfies the single crossing property, and the information structure is ordered according to the monotone likelihood ratio, then the comparative statics exhibit monotonicity with respect to signals. We consider the restrictions placed by this model of signal processing on state conditional stochastic choice data. In particular, we show that this model rationalizes a state conditional stochastic choice dataset if and only if the dataset itself is ordered according to the monotone likelihood ratio. A straightforward application of the main result shows the conditions under which the analyst can infer when one DM is more informed than the other.
@@
The single crossing property is an ordinal generalization of the increasing differences property [^4] A function $u$ exhibits increasing differences in $(x,\theta)$ if, $x^{\prime}>x$ and $\theta^{\prime}>\theta$ imply $u(x^{\prime},\theta^{\prime})-u(x,\theta^{\prime})>u(x^{\prime},\theta)-u(x,\theta)$. Clearly, this is a cardinal property. See Topkis, 1998.. Since it implies monotone comparative statics, and due to its straightforward economic interpretation resting on strategic complementarity, the SCP has turned out to be an extremely important tool for applications - both in applied theory as well as in the empirical literature. For instance, in mechanism design, single crossing property allows us to pin down optimal transfers. In applied work, an intimately related property has been used to solve certain optimal taxation problems (see Mirrlees, 1971). In IO entry models, single crossing property is often invoked and facilitates equilibrium analysis. In labor economics, the single crossing property is closely related to the so called Le Chatelier principle which helps characterize comparative static results (see Dekel, Quah and Sinander, 2022 for a recent work). The single crossing property also has some nice implications for preference aggregation, which proves to be quite useful in applied work (see, for instance Gans and Smart, 1996).
@@
In this paper we ask, what restrictions does the single crossing property impose on certain kinds of observable data? Often, available data - especially experimental data - may take a stochastic form. The analyst may observe only some distribution(s) over available actions and may not possess any knowledge of either the DM’s utility function or her private information. It is therefore important to understand whether the behavior of a Bayes rational agent imposes any restrictions on the stochastic data observed by the analyst. Caplin and Martin, 2015 show that rationalizability by a Bayes rational agent with access to some private information is equivalent to a single condition. Here, we ask what if the single crossing and MLR-ordered information structure impose any additional restrictions on top of those identified by Caplin and Martin, 2015. The importance of this question is illustrated by the aforementioned examples. Since the SCP has proven to be such an important property, it seems worthwhile to see if it imposes meaningful testable restrictions on observable data.
@@
Athey’s results suggest that the stochastic data should exhibit certain kind of monotonicity. If $(i)$ the DM’s utility function is single crossing; and $(ii)$ private information structure is MLR-ordered, then we know from Athey, 2002 that $(iii)$ the comparative statics are monotone. How does (i)(iii)(i)-(iii) affect the data? In our setup, neither the utility function nor the private information structure is observed. Only a set of conditional distributions over actions are observed. If $(ii)$ holds, then higher states lead to higher signal realizations more often - or with higher likelihood. If $(i)$ holds as well, then $(iii)$ is true, and so higher signals would lead directly to higher actions. One can expect this to be exhibited in data: since higher signals are more likely, and they lead to higher actions, we would expect higher actions to be more likely at higher states. This monotonicity translates into the data itself being MLR-ordered - that is, higher actions have higher relative probabilities at higher states. Proposition 13 shows that this intuition holds for the binary case. In fact, we establish something much stronger for the binary case. We show that this condition is not only both necessary and sufficient for rationalization by an SCP utility function, but such data can only be rationalized by a utility function satisfying the single crossing property. Furthermore, the main result of the paper shows that for any finite number of states, the MLR condition is both necessary and sufficient for rationalizability by a single crossing utility and an MLR ordered information structure.
@@
The main result suggests an interesting application. Because an MLR ordered dataset implies that that analyst may tentatively accept the hypothesis of a DM with single crossing utility and MLR ordered information. The Lehmann order ranks information structures based on their information content. In particular, one MLR-ordered information structure is Lehmann higher than the other if and only if, for every single crossing utility function, it yields a higher ex-ante expected utility. It is shown that, if two datasets are Lehmann ordered, then so are the information structures that rationalize them. This would allow the analyst to conclude that the DM to which the former dataset corresponds is in some sense revealed to be more informed. That is, the result allows an analyst to perform inference on how informed DMs just from choice data.
@@
Related literature. Our paper tries to answer a question similar to ones asked within the literature on revealed preference theory, although the dataset we consider is quite different from finite data typically considered in classical revealed preference theory. Green and Osband, 1991, Green and Srivastava, 1986, Echenique and Saito, 2015, Bayer et al., 2013 are prominent examples of classical revealed preference theory for expected utility. However, this set of papers considers a much richer setting. They focus primarily on consumption or asset demand, while none explicitly feature information processing or consider the single crossing property. Caplin and Martin, 2015, Doval and Eilat, 2023, Rehbeck, 2023 consider stochastic datasets and consider rationalizability by a Bayesian DM. The latter two consider rationalizability when the analyst only observes marginal distribution(s) over actions, but knows the DM’s utility function. Our paper can thus be thought of as imposing particular shape restrictions in the setting of Caplin and Martin, 2015. Lazzati, Quah and Shirai, 2018 consider similar shape restrictions, but they prove a rationalizability theorem with finite data, where the variation in data comes from “budgets” of actions. With finite data, they show that single crossing is indistinguishable from a weaker property developed by Quah and Strulovici, 2009 - called interval dominance property. They also derive similar results for cross-sectional datasets and Bayesian games.
@@
Apesteguia, Ballester and Lu, 2017 posits a single crossing random utility model, dubbed SCRUM. The primary difference between that paper and ours is contextual: they deal with stochastic choice - there is a set of preferences with a probability measure defined on them that needs to be recovered. Here, the preferences are not stochastic - only the data is - and one is interested in the existence of some preference and information structure consistent with the given data. Put another way, while the structure of the dataset available to the analyst is somewhat similar across these two groups of papers, the postulated data generating process is different.
@@
The next section presents the model, characterizes the dataset available to the analyst, and defines the corresponding notion of rationalizability. Section 2 presents the main result for the binary case - where the single crossing is indistinguishable from Bayesian expected utility maximization. Section 3 presents an example, and the two sections that follow establish necessity and sufficiency, respectively, for the general case. Section 6 shows how the main result can be used to perform inference on relative “informedness” of DMs, and the final section presents concluding discussion.
@@ 1.1 Preliminaries
Let $\Theta$ be a finite set of states and let $A$ be a finite set of actions, with $|A|\geq|\Theta|$. Typical elements of these sets are denoted by $\theta$ and $a$, respectively. Suppose that both $\Theta$ and $A$ are (strictly) ordered according to some exogenously given binary relations $_{\Theta},>_{A}$. [^5] In particular, $_{\Theta},>_{A}$ are assumed to be strict linear orders. That is, they are complete, transitive, and contain no “equalities”. While the results on which this paper builds are given for partial orders, we restrict attention to linear orders for brevity. Presence of unordered alternatives/states/signals does not affect the results reported here. To avoid notational clutter, in what follows, the subscripts are suppressed and the symbol >> is used to denote both the orderings. A decision maker (DM) wants to choose an action $a\in A$ without observing the state so as to maximize her expectation of her utility function $u:A\times\Theta\rightarrow\mathbb{R}$ over states.
Let $S$ be some set of signal realizations, with $|S|\geq|\Theta|$. Let $S$ be ordered by some fixed binary relation $_{S}$ as well. An information structure is a collection $\mathcal{I}\equiv\{\mu(\cdot|\theta)\}_{\theta\in\Theta}\subset\Delta(S)$. That is, each state induces a distribution over signal realizations. Let $\mu_{0}$ be the prior distribution over $\Theta$ [^6] To avoid unnecessary complications in the exposition, we assume that $\mu_{0}(\theta)>0$ for each $\theta\in\Theta$. This is without loss of generality since, if any state $\theta$ has a 0 probability, we can just drop that state and restrict attention to the set of states $\Theta/\{\theta\}$ instead. For revealed preference exercises, any such state is irrelevant.. The decision maker is said to be Bayesian if she obtains a signal realization $s\in S$, forms a belief $\mu(\cdot|s)\in\Delta(\Theta)$ using Bayes rule, and chooses an action $a^{*}(s)\in\arg\max_{A}\mathbb{E}_{\mu(\cdot|s)}[u(a,\theta)]$.
@@ 1.2 Data and rationalizability
Note, however, that it is possible that the DM optimally mixes between different actions. To that end, rationalizing the observed data also requires a choice rule $C:\Delta(\Theta)\rightarrow\Delta(A)$, mapping posterior beliefs into distributions over actions. Let $C(a|\gamma)$ denote the probability on action $a$ placed by the choice rule when the induced posterior is $\gamma$. Thus, an MBEU-DM is characterized by a tuple $\langle u,\mathcal{I},C\rangle$, where $u$ is a strict single crossing utility function, $\mathcal{I}$ is an MLR-ordered information structure, and $C$ is a choice rule.
We say that the data $\mathcal{D}$ is rationalized by an MBEU-DM - represented by $\langle u,\mathcal{I},C\rangle$ - if the distribution over actions induced by the optimal behavior of the DM characterized by the tuple is exactly $\bm{q}$. The following definition makes this notion of rationalizability precise,
@@
[…]
@@
The third condition demands that the observed data must in fact be induced by the choice behavior of the DM, while the fourth condition demands that the choice behavior be optimal. In certain degenerate cases with multiple optimal choices, it is possible that the choice rule does not select an otherwise optimal action. This may result in misleading inference. To address this issue, we assume that each optimal action is taken with some positive probability.
Assumption 1.
If $C$ is a choice rule corresponding to a DM with utility function $u$, then for each $\gamma\in\text{supp}(\pi),C(a|\gamma)>0$ if and only if $a\in\arg\max\left\{\mathbb{E}_{\gamma}[u(a,\theta)]\right\}$
While strong, the assumption is in line with the present context of complete observability. We now state the main result of the paper:
Theorem 1.
Consider a dataset $\mathcal{D}=\langle A,\Theta,\mu_{0},\bm{q}\rangle$.
  1. 1.
    If Assumption 1 holds and $\mathcal{D}$ corresponds to an MBEU-DM, then ${q(\cdot|\theta)}_{\theta\in\Theta}$ is MLR-ordered.
  2. 2.
    If ${q(\cdot|\theta)}_{\theta\in\Theta}$ is MLR-ordered, then $\mathcal{D}$ is MBEU-rationalizable.
2 Binary-state, binary-action case
In the two-states case, the single crossing property does not add any additional restrictions on top of those implied by expected utility maximization. Let there be two states and two actions so that $\Theta=\{\theta_{1},\theta_{2}\}$, $A=\{a_{1},a_{2}\}$, $a_{1}<a_{2}$ and $\theta_{1}<\theta_{2}$, and let the prior be $\mu_{0}=Pr(\theta=\theta_{1})$. The utility function is characterized by four numbers $u(a_{i},\theta_{j}),i,j=1,2$. The SCP is violated in this case if (and only if) $u(a_{1},\theta_{1})\leq u(a_{2},\theta_{1})$ and $u(a_{1},\theta_{2})>u(a_{2},\theta_{2})$. Denote the points in the dataset $q(a_{i}|\theta_{j})$ as $q^{i}_{j}$ - so that subscripts correspond to states and the superscripts correspond to actions. Similarly, denote the utility $u(a_{i},\theta_{j})$ by $u^{i}_{j}$. Note that any dataset in this case is characterized by two numbers: $q(a_{1}|\theta_{1})\equiv q^{1}_{1},q(a_{1}|\theta_{2})\equiv q^{1}_{2}$. From these, we can compute $q^{2}_{1}=1-q^{1}_{1}$ and $q^{2}_{2}=1-q_{2}^{1}$. The MBEU hypothesis then has a straightforward test as demonstrated in the following lemma.
@@ 2 Binary-state, binary-action case
Thus, in the binary case, the MLR property alone is both necessary and sufficient for there to exist an MBEU-rationalization. The above result suggests a straightforward test of the SCP. Indeed, if the data is not MLR-ordered, then no BEU-rationalization of the data can satisfy the single crossing property. On the contrary, if the data is MLR-ordered, then there exists a BEU-rationalization that satisfies the SCP. The proof also exhibits something much stronger. It turns out that, in the binary case, an MLR-ordered dataset can be rationalized only by an MBEU-DM. That is, if there is an information structure and a utility function that rationalizes the data, then the information structure must be MLR-ordered and the utility function must be single crossing. This result is stated below as a corollary.
@@
The result says that, in 2 action, 2 state case, Bayesian expected utility hypothesis is indistinguishable from single crossing property. Put another way, the single crossing property puts no additional restricitions on the expected utility theory. These restrictions on From an empirical perspective, this result is crucial, since many econometric application of single crossing property involve binary states (for instance, entry games). The next sections establish that Proposition 1 holds in the general case as well.
3 An Example
Let there be three equi-likely states $\theta_{1}<\theta_{2}<\theta_{3}$ and three available actions $a_{1}<a_{2}<a_{3}$. Consider the utility function and an information structure with three signal realizations $l<m<h$ given in tables 1 and 2 respectively.
Table 1: $u(a,\theta)$
$a_{1}$ $a_{2}$ $a_{3}$
$\theta_{1}$ 5 -1 2
$\theta_{2}$ -1 5 -1
$\theta_{3}$ 0 -1 5
Table 2: $\mathcal{I}\equiv\{\mu(s|\theta)\}$
$l$ $m$ $h$
$\theta_{1}$ 2/3 1/6 1/6
$\theta_{2}$ 1/4 1/2 1/4
$\theta_{3}$ 1/6 1/6 2/3
The function $u$ exhibits two violations of the single crossing property in $(a,\theta)$: (1) $u(a_{2},\theta_{1})<u(a_{3},\theta_{1})$, $u(a_{2},\theta_{2})>u(a_{3},\theta_{2})$, and (2) $u(a_{1},\theta_{2})<u(a_{2},\theta_{2})$, $u(a_{1},\theta_{3})>u(a_{2},\theta_{3})$.
Given the information structure, the distributions induced by the signals $l,m,$ and $h$ are $(8/13,3/13,2/13),(2/10,6/10,2/10)$, and $(2/13,3/13,8/13)$ respectively. These distributions are, in turn, induced with probabilities $13/36,10/36,13/36$ respectively. The optimal actions are $a^{*}(l)=a_{1}$, $a^{*}(m)=a_{2}$, $a^{*}(h)=a_{3}$. Thus, the induced state-conditional distributions over actions are,
$\mu(\cdot|\theta_{1})$ $=(2/3,1/6,1/6)$
$\mu(\cdot|\theta_{2})$ $=(1/4,1/2,1/4)$ (2)
$\mu(\cdot|\theta_{3})$ $=(1/6,1/6,2/3)$
In the present context, the stochastic data available to the analyst takes the form of equation 2. The question this paper poses is, under what conditions on $\{\mu(\cdot|\theta)\}_{\theta_{\in}\Theta}$, does there exist an information structure, an SCP utility function, and a choice rule [^14] Indeed, the choice rule may be degenerate, mapping posteriors into Dirac measures over actions. consistent with the former in the sense of definition 4.
Note, however, that the above utility function may be changed to a best-response equivalent utility function which has no violations of the SCP. One such utility function is depicted in table 3 below.
Table 3: A $u(a,\theta)$ satisfying the SCP
$l$ $m$ $h$
$\theta_{1}$ 5 3 1
$\theta_{2}$ 3 5 -1
$\theta_{3}$ -1 1 10